How much energy is produced by a reaction of 33.5 g of iron with excess oxygen: 3Fe+2O2---->Fe3O4+1120kJ
1
Expert's answer
2020-05-19T08:57:59-0400

Moles of iron in "33.5\ g=\frac{33.5}{55.84}=0.6"

"3Fe+2O_2\to Fe_3O_4+1120\ KJ"

"3" moles of Iron produce "1120\ KJ" of energy.

"0.6" moles will produce "\frac{0.6}{3}\times 1120=224\ KJ" of energy.